OpenStudy (star_girl):

Please help! This question is a simple calculating interest question. I've tried it many times and am always a little bit off of the answer. I'm not sure what I am doing wrong.

OpenStudy (star_girl):

George charged a motorized bicycle on his credit card for $2,000 on June 12. His balance before that purchase on that credit card was $500. His credit card company calculates interest using the ADB including new purchases method. The APR on George’s credit card is 15% APR. Hint: the credit card rounded the monthly interest rate up to .013%. George’s finance charge for this month was $22.97. Please show how to calculate this interest.

OpenStudy (mathmate):

There's something missing or assumed in the question: the billing cycle. Normally credit cards have a spread out billing cycle so as not to have a rushed billing period at the end of the month. For lack of information, we will ASSUME that the billing cycle is the calendar month of June, namely 30 days. If the monthly interest is 0.013 (instead of 0.013%) exactly, then the accrued interest is 0.013((19 days with balance of 2500)/30+(11 days with balance of 500)/30) =0.013(2500(19/30)+500(11/30)) =22.97 Note: 0.013% monthly interest corresponds to an annual interest of 0.156%, which is more than 100 times lower than the current interest rates of 19 to 25% per annum.

OpenStudy (star_girl):

Thank you so much @mathmate!!!!

OpenStudy (mathmate):

You're welcome! :)